# Think Smart, Act SMARL! Analyzing Probabilistic Logic Driven Safety in Multi-Agent Reinforcement Learning

### v1

@@
Abstract.
An important challenge for enabling the deployment of reinforcement learning (RL) algorithms in the real world is safety. This has resulted in the recent research field of Safe RL, which aims to learn optimal policies that are safe. One successful approach in that direction is probabilistic logic shields (PLS), a model-based Safe RL technique that uses formal specifications based on probabilistic logic programming, constraining an agent’s policy to comply with those specifications in a probabilistic sense. However, safety is inherently a multi-agent concept, since real-world environments often involve multiple agents interacting simultaneously, leading to a complex system which is hard to control. Moreover, safe multi-agent RL (Safe MARL) is still underexplored. In order to address this gap, in this paper we ($i$) introduce Shielded MARL (SMARL) by extending PLS to MARL – in particular, we introduce Probabilistic Logic Temporal Difference Learning (PLTD) to enable shielded independent Q-learning (SIQL), and introduce shielded independent PPO (SIPPO) using probabilistic logic policy gradients; ($ii$) show its positive effect and use as an equilibrium selection mechanism in various game-theoretic environments including two-player simultaneous games, extensive-form games, stochastic games, and some grid-world extensions in terms of safety, cooperation, and alignment with normative behaviors; and ($iii$) look into the asymmetric case where only one agent is shielded, and show that the shielded agent has a significant influence on the unshielded one, providing further evidence of SMARL’s ability to enhance safety and cooperation in diverse multi-agent environments.
Key words and phrases:
Multi-Agent Reinforcement Learning, Neurosymbolic AI, AI Safety, Game Theory, Probabilistic Logic Programming
1. Introduction
@@
Recent years have witnessed significant progress in multi-agent reinforcement learning (MARL), with sophisticated algorithms tackling increasingly complex problems. These advancements have led to remarkable progress in various domains, including autonomous vehicles where fleets of self-driving vehicles are expected to safely coordinate to optimize traffic flow and reduce congestion (Shalev-Shwartz et al., 2016), robotic swarms where multiple robots collaborate to complete tasks more efficiently (Brambilla et al., 2013), trading agents that can operate in tandem to maximize returns and manage risk more effectively (Ganesh et al., 2019), energy grid management where multiple agents control different aspects of the power grid to balance supply and demand efficiently (van der Sar et al., 2023), and healthcare systems where agents coordinate to optimize scheduling and resource allocation for patience and hospitals (Riek, 2017), just to name a few. The ultimate success of RL algorithms in this diverse collection of domains and the deployment in the real-world, however, demands overcoming a single yet difficult key challenge: safety.
@@
This has resulted in a recent push in the research field of Safe RL which aims to learn optimal policies that are safe, including Hunt et al. (2021); Jansen et al. (2020); Den Hengst et al. (2022) among others. A common property of these approaches is that they use linear temporal logic (LTL) to represent safety constraints. One recent successful proposal that also uses a formal language to represent safety constraint is probabilistic logic shields (PLS) (Yang et al., 2023) whose semantics is based on probabilistic logic programming. PLS constrains agent’s policy to comply with formal specifications probabilistically, and comes with various advantages over other techniques: ($i$) it is a probabilistic shield that is applied at the policy level instead of individual actions capturing the probabilistic evaluation of safety in contrast to deterministic rejection based shields Hunt et al. (2021); Jansen et al. (2020); ($ii$) PLS is less demanding in terms of requiring knowledge of the underlying MDP as opposed to Hunt et al. (2021); Jansen et al. (2020); Carr et al. (2023); ($iii$) being end-to-end differentiable, it integrates logical semantics to deep learning architectures seamlessly; and ($iv$) provides safety guarantees for (single) agents.
@@
However, safety is inherently a multi-agent concept, since real-world environments often involve multiple agents interacting simultaneously, leading to a complex system which is hard to control. While there are a few safe MARL approaches trying to tackle safety in the multi-agent setting Gu et al. (2023); ElSayed-Aly et al. (2021); Melcer et al. (2022), to the best of our knowledge, PLS has not been adopted, extended or studied in this way. In order to address this gap, in this paper:
  1. (1)
    We introduce Shielded MARL (SMARL) by extending PLS to MARL – in particular, we introduce Probabilistic Logic Temporal Difference Learning (PLTD) to enable shielded independent Q-learning (SIQL), and introduce shielded independent PPO (SIPPO) using probabilistic logic policy gradients;
  2. (2)
@@
    We show that PLS can be used as a equilibrium selection mechanism in various 2-player game-theoretic environments including a two-player simultaneous game (Stag-Hunt), an extensive-form games (Centipede), a stochastic game (Extended Public Goods Games), and a grid-world extension (Markov Stag-Hunt) in terms of safety and cooperation. Moreover, we investigate the impact of smaller (weak) and larger (strong) shields;
  3. (3)
    We investigate the influence of shielded agents over unshielded ones. Results suggest that the agent equipped with the shield can guide the other agent’s behavior significantly, suggesting further evidence of SMARL’s ability to enhance safety and cooperation in diverse multi-agent environments.
1.1. Related Work
In the domain of safe MARL, a few but significant strides have been made recently in developing frameworks that ensure safety during learning and execution. A relatively recent one, Gu et al. (2023), propose constrained Markov game and a constrained multi-agent policy gradient optimization algorithms (MACPO), however without the use of any formal language semantics. More related works include ElSayed-Aly et al. (2021) which introduce a framework that integrates shielding mechanisms to dynamically enforce safety constraints. Extending this, Melcer et al. (2022) explore decentralized shielding, i.e. allowing each agent to have its own shield, reducing computational overhead and enhancing scalability. Both of these approaches, use LTL as formal language, and the higher requirement to know the underlying MDP – this is in contrast to our use of a probabilistic logic framework (ProbLog) and weaker demand of having access to the MDP (only the safety-related parts of the states), respectively.
2. Preliminaries
In this section, we provide a brief overview of the formal basis for relevant game theoretic preliminaries, multi-agent reinforcement learning and probabilistic logic shielding following Yang et al. (2023).
2.1. Game Theory
@@
A normal-form game (NFG) is a tuple $\langle N,\mathbf{A},\mathbf{u}\rangle$ where $N=\{1,...,n\}$ is a finite set of agents (or players), $\mathbf{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ is a finite set of action (also called pure strategy) profiles $\mathbf{a}=(a_{1},...a_{n})$ with $\mathcal{A}_{i}$ being the set of player $i$’s actions, and $\mathbf{u}=(u_{1},...,u_{n})$ is a profile of utility functions $u_{i}:\mathbf{A}\rightarrow\mathbb{R}$. A strategy profile $\mathbf{s}=(s_{1},...s_{n})$ is called a Nash equilibrium if for each player $k\in N$, the strategy $s_{k}$ is a best response to the strategies of all the other players $s_{i\in N\setminus\{k\}}$. Although we provided the more general $n$-player definition, in the experiments and the remainder of the paper, we shall only consider the 2-player variants. A generic payoff matrix (another representation of the utility function) for NFGs is in Figure 1(a).
@@
(a) Generic 2-player payoff matrix (e.g. Stag-Hunt)
(b) Expected EPGG utilities with endowments $c_{1}$ and $c_{2}$
(c) Centipede Game instance
(d) Markov Stag-Hunt (Peysakhovich and Lerer, 2018) environment instance
Figure 1. Representations of the games.
2.2. Safe MARL
2.2.1. Constrained MDPs
RL was developed to enable an single agent to learn a policy $\pi$ maximizing the sum of discounted reward signals collected in a stochastic and generally stationary environment (Nowé et al., 2012). The most common formalism for RL utilizes the notion of Markov Decision Processes (MDPs) (van Otterlo and Wiering, 2012; Sutton and Barto, 2018). An MDP is defined by a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a transition function $\mathcal{T}$, and a reward function $\mathcal{R}$. The objective in an MDP is to find a policy that maximizes the expected cumulative reward over time. However, this does not inherently account for other important considerations such as safety or costs associated with actions that might be undesirable in real-world scenarios. Constrained MDPs (CMDPs) are an extension of MDPs that include constraints, typically related to safety or resource usage (Altman, 2021). These additional considerations are handled through constraints that limit the policies to those that not only maximize rewards but also satisfy certain conditions, such as remaining within a specified cost or risk limit. Formally, a CMDP has a set of cost functions $\mathcal{C}_{i}$ for $i\in\{1,...,m\}$, each associated with its own threshold $d_{i}$. The goal is to maximize the cumulative reward while ensuring that the expected cumulative cost for each $\mathcal{C}_{i}$ does not exceed its respective threshold $d_{i}$.
2.2.2. MARL
MARL extends traditional RL paradigms such as PPO (Schulman et al., 2017) and Q-Learning (Mnih et al., 2013) to settings involving multiple agents that interact within a shared environment (respectively, De Witt et al., 2020; Tan, 1993). Each agent optimizes its own individual reward, but the presence of other co-learning agents introduces challenges such as non-stationarity, as the environment changes with the actions of the other agents (Albrecht et al., 2023). MARL finds applications in areas like autonomous driving (Shalev-Shwartz et al., 2016), financial trading (Ganesh et al., 2019) and energy grid management (van der Sar et al., 2023), where some combination of coordination, cooperation and competition is often required. Ensuring safety and collaboration among agents in MARL is a crucial issue, as unsafe actions can lead to significant consequences in high-stakes domains.
2.3. Probabilistic Logic Shielding
We summarize Yang et al. (2023) in describing Probabilistic Logic Shielding (PLS), a probabilistic method for safety in RL. Unlike traditional rejection-based shields that fully block unsafe actions and limit exploration (Alshiekh et al., 2018), PLS adjusts the probability of unsafe actions proportional to their risk, allowing them to occasionally occur during training. This enables the agent to learn from unsafe actions and improve its safety policy over time, particularly in environments with soft safety constraints.
Assume a probabilistic model $P$ that is aware of the probability $P(\texttt{safe}|a,s)$ of the safety of taking a given action $a$ for a given state $s$. $P$ may not necessarily represent the whole environment or transition model (i.e. it need not be a full representation of the underlying MDP), but only encode relevant knowledge about the safety of certain states and how they relate to the current policy $\pi$. The safety of $\pi$ is the sum of the disjoint probabilities of the safeties of taking each action:
@@ 2.3. Probabilistic Logic Shielding
In order to compute a safer policy, we can marginalize actions out with respect to $\pi$. This leads us to the notion of probabilistic shielding: formally, given a base policy $\pi$ and a probabilistic safety model $P(\mathtt{safe}|s,a)$, the shielded policy is
@@
Yang et al. (2023) choose to implement the model $P$ in ProbLog (De Raedt et al., 2007), as opposed to other representations, for a number of reasons, including that it is easily differentiable and allows easy modeling and planning. This means our safety constraints are to be specified symbolically through probabilistic logical predicates and relations.
@@
Within this paper, a shield is synonymous with a ProbLog program $\mathcal{T}$ that takes in two sets of inputs: the first is the action distribution from a policy $\pi(a|s)$ ($\forall a\in\mathcal{A}$). The second is certain information that may come from the agent or the environment, each one being input via a ‘sensor’. These are not necessarily the same as the information that is input to the reinforcement learning model. Using such a ProbLog program to perform probabilistic shielding is called probabilistic logic shielding. The appropriate ProbLog program $\mathcal{T}$ induces a probabilistic measure $\mathbf{P}_{\mathcal{T}}$, and we can (conditionally) query the program to give us various information including the action safety (for each action $a$) as $P(\texttt{safe}|s,a)=\mathbf{P}_{\mathcal{T}}(\texttt{safe}|a)$, the safety of the whole policy under $s$ as $P_{\pi}(\texttt{safe}|s)=\mathbf{P}_{\mathcal{T}}(\texttt{safe})$ and the safe/shielded policy $\pi^{+}$ as $P_{\pi}(a|s,\texttt{safe})=\mathbf{P}_{\mathcal{T}}(a|\texttt{safe})$.
2.4. ProbLog Shields

### v2

# Analyzing Probabilistic Logic Shields for Multi-Agent Reinforcement Learning

@@
Abstract
Safe reinforcement learning (RL) is crucial for real-world applications, and multi-agent interactions introduce additional safety challenges. While Probabilistic Logic Shields (PLS) has been a powerful proposal to enforce safety in single-agent RL, their generalizability to multi-agent settings remains unexplored. In this paper, we address this gap by conducting extensive analyses of PLS within decentralized, multi-agent environments, and in doing so, propose Shielded Multi-Agent Reinforcement Learning (SMARL) as a general framework for steering MARL towards norm-compliant outcomes. Our key contributions are: (1) a novel Probabilistic Logic Temporal Difference (PLTD) update for shielded, independent Q-learning, which incorporates probabilistic constraints directly into the value update process; (2) a probabilistic logic policy gradient method for shielded PPO with formal safety guarantees for MARL; and (3) comprehensive evaluation across symmetric and asymmetrically shielded $n$-player game-theoretic benchmarks, demonstrating fewer constraint violations and significantly better cooperation under normative constraints. These results position SMARL as an effective mechanism for equilibrium selection, paving the way toward safer, socially aligned multi-agent systems.
[…]
4833
1 Introduction
@@
Recent years have witnessed significant progress in multi-agent reinforcement learning (MARL), with sophisticated algorithms tackling increasingly complex problems in various domains including autonomous vehicles [28], distributed robotics [5], algorithmic trading [13], energy grid management [34] and healthcare [26]. The ultimate success of RL in this diverse collection of domains and the deployment in the real-world, however, demands overcoming a difficult key challenge: safety.
This has naturally resulted in the research direction of Safe RL which aims to learn optimal policies that are, by some measure, ‘safe’. Gu et al. [15] present an up-to-date overview of the field. A number of proposals focus on the application of formal methods [16, 11], within which the notion of shielding is used, a technique that is inspired by formal verification through temporal logic specifications to avoid unsafe actions during the agent’s learning process [4, 1, 17, 7].
@@
One recent proposal that uses shielding to represent safety constraints is probabilistic logic shields [PLS, 39] whose semantics is based on probabilistic logic (PL) programming [9]. PLS constrains an agent’s policy to comply with formal specifications probabilistically. The specification of constraints (the shield), is defined within the exploration and learning pipeline. PLS comes with various advantages over other techniques: ($i$) the shields are applied at the policy level instead of individual actions capturing the probabilistic evaluation of safety (in expectation) in contrast to deterministic shields that are based on hard-rejection [16, 17]; ($ii$) PLS is less demanding in terms of requiring knowledge of the underlying MDP compared to previous methods [e.g., 16, 17, 7]; ($iii$) being end-to-end differentiable, it integrates logical semantics with deep learning architectures seamlessly; and ($iv$) it provides safety guarantees within single-agent settings.
Figure 1: Shielded MARL (SMARL) interaction schematic. At time $t$, each agent $p_{i}$ passes their respective policy $\pi_{i}^{t}(s_{i}^{t})$ and a safety-specific state $s^{t}_{i}$ to their shield $\mathcal{T}_{i}$, resulting in a safe policy $\pi_{i}^{+}(s_{i}^{t})$ from which an action $a^{t}_{i}$ is sampled and returned to the environment.
@@
However, safety is inherently a multi-agent concept, as real-world environments often involve multiple agents interacting simultaneously, leading to hard-to-control complex systems. While there exist a few approaches tackling safety in multi-agent settings [14, 12, 20, 29], to the best of our knowledge, PLS has not been adopted, extended or analyzed in the context of MARL. In this paper, we address this gap with the following contributions:
  1. 1.
    We present a framework that extends shielding in RL to multi-agent settings, named Shielded MARL (SMARL). We provide a theoretical guarantee that SMARL with PLS produces safer joint policies than unshielded counterparts. Using decentralized techniques, we introduce (a) Probabilistic Logic Temporal Difference Learning (PLTD), with convergence guarantees, allowing for shielded independent Q-learning (SIQL), and (b) shielded independent PPO (SIPPO) using probabilistic logic policy gradients;
  2. 2.
@@
    We show that PLS can be used as an equilibrium selection mechanism, a key challenge in MARL, under various game-theoretic settings, such as coordination, spatio-temporal coordination, social dilemmas, cooperation under uncertainty, and mixed-motive problems. We provide empirical evidence across various $n$-player environments including an extensive-form game (Centipede), a stochastic game (Extended Public Goods Game), a simultaneous game (Stag-Hunt), and its grid-world extension (Markov Stag-Hunt). Moreover, we investigate the impact of smaller (weak, less specific) and larger (strong, more specific) shields;
  3. 3.
    We investigate asymmetric shielding, i.e. the ability of shielded agents to influence unshielded peers. Results show that partial shielding can significantly enhance safety, highlighting SMARL’s effectiveness in both cooperative and non-cooperative settings.
(a) Payoff matrix for Stag-Hunt
(b) Expected EPGG utilities
(c) Centipede instance
(d) Markov Stag-Hunt [24]
Figure 2: Representations of the games that are experimented within this paper. For details please refer to Section 2.2.
2 Preliminaries
We start by specifing what we mean by RL safety, provide background knowledge on the game-theoretic environments we use, and finally a quick introduction to PLS.
2.1 Safety Definition
We expand the interpretation of ‘safety’ in RL beyond that of works like ElSayed-Aly et al. [12] and Yang et al. [39] (i.e. informally, the agent should not perform actions that may lead to something ‘harmful’), and more towards an alternative conception provided in Gu et al. [15, pg. 3], namely, they act, reason, and generalize obeying human desire. We formalize this in the vein of Alshiekh et al. [1]:
Definition 2.1.
Formally-Constrained Safe RL is the process of learning an optimal policy $\pi^{*}$ while maximally satisfying a set of formal specifications $\mathcal{T}$ during learning and deployment.
Such formal specifications may represent any constraints including arbitrary temporal logic formulae or, in multi-agent settings, social norms or equilibrium conditions, which may not necessarily be ‘harmful’, but may instead ascribe normative or even desirable behavior. In our case, $\mathcal{T}$ is encoded using a PL program.
2.2 Game Theory
@@
A normal-form game (NFG) is a tuple $\langle N,\mathbf{A},\mathbf{u}\rangle$ where $N=\{1,...,n\}$ is a finite set of agents (or players), $\mathbf{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ is a finite set of action profiles $\mathbf{a}=(a_{1},...a_{n})$ with $\mathcal{A}_{i}$ being the set of player $i$’s actions, and $\mathbf{u}=(u_{1},...,u_{n})$ is a profile of utility functions $u_{i}:\mathbf{A}\rightarrow\mathbb{R}$. A strategy profile $\mathbf{s}=(s_{1},...s_{n})$ is called a Nash equilibrium if for each player $k\in N$, the strategy $s_{k}$ is a best response to the strategies of all the other players $s_{i\in N\setminus\{k\}}$.
@@
2.3 Probabilistic Logic Shielding
We summarize the approach of Yang et al. [39], known as Probabilistic Logic Shielding (PLS), a method for incorporating probabilistic safety in RL, and refer the reader to the original work for a more detailed treatment. Unlike traditional rejection-based shields that fully block unsafe actions and limit exploration [1], PLS adjusts the probability of unsafe actions proportional to their risk, allowing them to occasionally occur during training. This lets agents learn from unsafe actions, especially under soft constraints.
Assume a model $P$ that assigns probabilities $P(\texttt{safe}|s,a)$ representing the likelihood that taking action $a$ in state $s$ is safe. $P$ may not necessarily represent the whole environment or transition model (i.e. it need not be a full representation of the underlying MDP), but only encode relevant knowledge about the safety of certain states and how they relate to an agent’s current policy $\pi$. The safety of $\pi$ is the sum of the disjoint probabilities of the safety of taking each action:
@@ 2.3 Probabilistic Logic Shielding
To obtain a policy that favors safer actions under $P$, we can marginalize actions out w.r.t. $\pi$. This leads us to the notion of probabilistic shielding: formally, given a base policy $\pi$ and a probabilistic safety model $P(\mathtt{safe}|s,a)$, the shielded policy is:
@@
Yang et al. [39] choose to implement the model $P$ in ProbLog [9] for a number of reasons, including that it is easily differentiable and allows easy modeling and planning. This means our safety constraints are to be specified symbolically through PL predicates and relations.
@@
Within PLS, a shield is synonymous with a ProbLog program $\mathcal{T}$ that receives two sets of inputs: the first is the action distribution from a policy $\pi(s)$, and the second is a set of probabilistic fact valuations encoding safety-related information that may come from the agent or the environment – this may differ from the state information used to compute the base policy $\pi$. The appropriate ProbLog program $\mathcal{T}$ induces a probabilistic measure $\mathbf{P}_{\mathcal{T}}$, and we can (conditionally) query the program to give us various information including the action safety (for each action $a$) as $P(\texttt{safe}|s,a)=\mathbf{P}_{\mathcal{T}}(\texttt{safe}|a)$, the safety of the whole policy under $s$ as $P_{\pi}(\texttt{safe}|s)=\mathbf{P}_{\mathcal{T}}(\texttt{safe})$ and the safe/shielded policy $\pi^{+}$ as $P_{\pi}(a|s,\texttt{safe})=\mathbf{P}_{\mathcal{T}}(a|\texttt{safe})$.
3 Probabilistic Logic Shields for Multi-Agent Reinforcement Learning
We present a general SMARL framework in Section 3.1, applying (probabilistic) shielding to multi-agent algorithms. Next, in Section 3.2, we introduce PLTD, which incorporates logical constraints directly into the TD-learning process. Finally, Section 3.3 outlines the encoding of safety constraints as PL programs.
3.1 General Framework Description
A schematic diagram of a parallel SMARL setup is in Figure 1. Each agent $p_{i}$ (indexed with $i\in\{1,...,n\}$) observes a possibly subjective state $s_{i}^{t}$ at time $t$, and sends both the computed policy $\pi_{i}(s^{t}_{i})$ and relevant safety-related information about $s_{i}^{t}$ to its shield $\mathcal{T}_{i}$. Each $\mathcal{T}_{i}$ is a ProbLog program that is used to compute a safe policy $\pi_{i}^{+}(s_{i}^{t})$ according to Eq. 2 from which an action $a_{i}^{t}$ is sampled. Finally, the joint action profile $\vec{a}^{t}$ is sent to the environment $Env$. After updating its state, $Env$ sends new observations $\mathbf{s}^{t+1}$ and rewards $\vec{\mathbf{r}}^{t+1}$ to the agents, and the cycle continues with $t\leftarrow t+1$ until a terminal state is reached. The agent-environment-cycle interaction scheme, used often in turn-based games [32] can likewise be extended naturally to the SMARL framework by having each agent interact with their own shield before sampling an action.
In Probabilistic Logic SMARL (PL-SMARL), at least one of the RL agents use PLS, optionally sharing policies and/or parameter updates. Two variations (using neural network approximators) of Independent Q-learning (IQL) [31] are introduced as Shielded IQL (SIQL), without parameter sharing, and Shielded Parameter-Sharing IQL (PSQL), sharing all parameters. Likewise, three variants of Independent PPO (IPPO) [10] are examined: Shielded IPPO (SIPPO), Critic-Sharing PPO (CSPPO) and Actor-Critic-Sharing PPO (ACSPPO).
3.1.1 Safety Guarantees
Following Definition 2.1, we define relative safety for SMARL:
Definition 3.1 (Relative SMARL Safety).
A joint policy $\vec{\pi}^{+}=\langle\pi_{1}^{+},\pi_{2}^{+},\ldots,\pi_{n}^{+}\rangle$ is defined to be at least as safe as another joint policy $\vec{\pi}=\langle\pi_{1},\pi_{2},\ldots,\pi_{n}\rangle$ if and only if for all agents $i\in N$ with respective shields $\mathcal{T}_{i}$, it is the case that $P_{\pi^{+}_{i}}(\texttt{safe}|s)\geq P_{\pi_{i}}(\texttt{safe}|s)$ for all reachable states $s\in S$; and is strictly safer if additionally for at least one $j\in N$ and $s\in S$, $P_{\pi^{+}_{j}}(\texttt{safe}|s)>P_{\pi_{j}}(\texttt{safe}|s)$.
Based on Definition 3.1, we attain the following proposition:
Proposition 3.2.
A PL-SMARL algorithm with agents $i\in N$ with respective policies $\pi_{i}^{+}$ and shields $\mathcal{T}_{i}$ has a joint shielded policy $\vec{\pi}^{+}$ at least as safe as their base joint policy $\vec{\pi}$, and strictly safer whenever any individual base policy $\pi_{j}$ violates $\mathcal{T}_{j}$.
A proof is provided in the supplementary materials. This is similar to the first case of ElSayed-Aly et al. [12]’s correctness proof of the safety of factored shields. However, unlike factored shields, PLS does not guarantee per-step safety; rather, it ensures that the expected safety under the shielded policy is higher than that of the base policy.
3.2 PL-Shielded Temporal Difference Learning (PLTD)
Yang et al. [39] define their shielded algorithm, PLPG, as an extension of policy gradient methods, specifically, PPO [27]. Thus, to use it with DQN [21], SARSA [41] (or other TD-learning algorithms), we begin by formalizing how a ProbLog shield is used in conjunction with TD-learning. An important consideration for extending PLS to TD-learning is whether the algorithm is on-policy or off-policy. Yang et al. [39] note that for off-policy algorithms to converge, both the exploration and learned policies must cover the same state-action space – this assumption must be handled with care, particularly in the design of exploration strategies. We propose both on- and off-policy algorithms that incorporate safety constraints into Q-learning.
3.2.1 Objective Function
PLPG includes a shielded policy gradient $[\nabla_{\theta}\log P_{\pi^{+}}(\mathtt{safe}|s)]$ and a safety gradient penalty $[-\log P_{\pi^{+}}(\mathtt{safe}|s)]$. Since TD-methods do not use policy gradients, we must rely on the latter to introduce information about safety constraints into the objective function.
Let $\mathcal{D}$ be the distribution of $d=\langle s^{t},a^{t},r^{t},s^{t+1},a^{t+1}\rangle$ tuples extracted from a history buffer of the agent interacting with the MDP and $s^{t}$, $a^{t}$, $r^{t}$ be the state, taken action, and reward at time $t$. Let the Q-value approximator $Q_{\theta}(s,a)$ be parameterized by some parameters $\theta$. The off-policy loss (Q-learning based) and on-policy loss (SARSA-based) are augmented using the aforementioned safety penalty term. We refer to these methods as Probabilistic Logic TD Learning (PLTD).
Definition 3.3 (Probabilistic Logic TD Learning).
The PLTD minimization objective is:
Under perfect sensor information, PLTD reduces to vanilla TD-learning within the (reachable) set of safe states, and thus in limited settings is guaranteed to converge to an optimal safe policy [33].
Proposition 3.4.
PLTD, i.e. Eq. 3 converges to an optimal safe policy given perfect safety information in all states in tabular and linear function approximation settings.
A proof of this is in the supplementary material. For deep neural Q-networks, global convergence is currently not known; in that regime we can adopt the usual stabilization tricks (target network, replay buffer, Double-DQN). While we observe empirical stability of PLTD, a complete theoretical analysis remains an open question.
3.3 Shield Construction using ProbLog